A printing machine can print 250 flyers per minute. A folding machine can fold the flyers at the rate of 5 per second. If 6 printing machines were being used to print the flyers, then _______ folding machines would be needed to keep up with the 6 printing machines.
Answer by richwmiller(17219)

5 per second =5*60=300 per minute
6*250=300x
6*25=30x
25=5x
x=5 folding machines